OpenStudy (anonymous):

Solve the equation. Show your work. 45 = 3b + 69

OpenStudy (anonymous):

@Notamathgenius help D:

OpenStudy (anonymous):

@yuriana

OpenStudy (anonymous):

@Geckoh2o

OpenStudy (notamathgenius):

Simplifying 45 = 3b + 69 Reorder the terms: 45 = 69 + 3b Solving 45 = 69 + 3b Solving for variable 'b'. Move all terms containing b to the left, all other terms to the right. Add '-3b' to each side of the equation. 45 + -3b = 69 + 3b + -3b Combine like terms: 3b + -3b = 0 45 + -3b = 69 + 0 45 + -3b = 69 Add '-45' to each side of the equation. 45 + -45 + -3b = 69 + -45 Combine like terms: 45 + -45 = 0 0 + -3b = 69 + -45 -3b = 69 + -45 Combine like terms: 69 + -45 = 24 -3b = 24 Divide each side by '-3'. b = -8 Simplifying b = -8

OpenStudy (anonymous):